OpenStudy (anonymous):

Which of the following is the slope between the points (–1, –5) and (7, –3)? A 4 B 1 over 4 C –1 over 4 D -4

OpenStudy (anonymous):

B) 1/4 Slope = (y2-y1)/x2-x1)

OpenStudy (ghazi):

m= (-3-(-5))/(7-(-1))=2/8

OpenStudy (anonymous):

THANK YOU both! :)

OpenStudy (ghazi):

stay happy :)